A Bibliography of Smithsonian Magazine: 2000–2009

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

07 August 2013 Version 2.02

Title word cross-reference

2000 [200].

A-changin [120]. Absurd [45]. Addictive [208]. Additional [14, 31, 49, 56, 89, 107, 124, 142, 159, 175, 194, 201]. Address [115]. Afloat [100]. Aim [9]. Air [155, 37]. Alert [132]. Alive [63]. Amateuring [90]. Amazing [105]. America [54, 118]. American [152, 210, 138, 7]. Andrew [192]. Anna [50]. Answered [62]. Antarctic [122]. Ants [10]. Any [30]. Anyone [8]. Appreciated [136]. Art [110, 83, 138, 78, 170, 46]. Artists [42]. Ashes [65]. Assembly [106, 112]. Attractions [183]. Author [105]. Avant [188]. Avant-Garde [188]. Away [64].

Baking [189]. Ball [212]. Barns [179].

Baseball [173]. Basque [70]. Bath [206]. Bats [72]. Battling [117]. Batty [84]. Be [158, 136]. Bears [10]. Beasts [32, 85]. Beauty [85]. Bedtime [30]. Bees [72]. Beirut [65]. Belongs [43]. Below [81]. Beneath [122]. Bernd [33]. Better [180]. Beyond [3, 55, 69, 78, 98, 133, 150, 184, 211]. Big [181, 6]. Biggest [64]. Birds [40, 72]. Boats [100]. Bonding [55]. Bone [69]. Bones [130]. Book [57, 176, 90, 15, 33, 143, 50, 210, 174, 30]. Books [186, 200, 55]. Booth [90]. Bound [79]. Bourron [152]. Bourron-Marlotte [152]. Bozeman [171]. Breck [11]. Bright [195]. Buffalo [21]. Building [180]. Bureau [77, 95]. Business [189]. Buy [172]. bye [13].

Calendar [16, 34, 51, 58, 91, 108, 125, 144, 160, 177, 196, 202]. Call [69]. Called [62]. Camping [96]. Can [71]. Can’t [158].

Canyon [57]. Capturing [118]. Careful [158]. Carroll [192]. Casting [24]. Cat [168]. Catalogs [193]. Celestial [77, 95]. Century [19]. changin [120]. Chardin [102]. Charting [99]. Child [151]. Children [186, 200]. Chillida [119]. Chips [5]. Circle [82]. Clifford [15]. Collection [73]. Collector [182]. Cometh [205]. Coming [48]. Commenced [94]. Comment [77, 132, 95, 39, 74, 167, 22]. Concert [96]. Control [193]. Controversy [97]. Corn [81]. Cotten [166]. Court [123]. Crafts [83]. Cutting [46]. Cutting-edge [46]. Cypress [213].

daisy [18]. Damadian [97]. Dancing [155]. Davis [62]. Day [123, 29]. Days [162, 158]. Dead [63, 68]. Deserts [24]. Designs [185]. Diamond [162]. Didn’t [64]. Diffusion [1]. Dinosaur [180]. Do [30, 72]. Does [18]. Dog [123]. Dogs [217]. Down [120]. Driven [54]. Drum [128]. Drums [98].

Earning [71]. Earns [36]. Easy [18]. Eat [10]. edge [46]. Editor [17, 35, 52, 59, 92, 109, 126, 145, 161, 178, 197, 203]. Eduardo [119]. Eels [28]. Empire [68]. End [20]. Even [72]. Every [123, 151]. Everyday [75]. Everywhere [121]. Evidence [114, 74]. Experience [78]. Explore [54].

Farewell [195]. Farm [120]. Faster [176]. Fatal [39]. Feast [15]. feeding [72]. Feet [81]. Fellow [158]. Feuds [39]. Fight [118]. Finding [121]. Fire [57, 9, 57]. Fishy [135]. Flag [116]. Flood [74]. Flower [50]. Fly [100]. Flying [84]. Foreign [132]. Forensics [39]. Forgotten [110, 87]. Founder [87]. Foxes [84]. Freedom [118]. French [210]. Fun [101]. Funky [101]. Fury [25]. Futile [39].

Gallery [150]. Garde [188]. Gavrilo [134]. Gem [73]. George [87]. Get [64]. Gift [204]. Girls [11]. Glacier [122]. Glass [93]. Gleick [176]. Glory [79]. Gold [44, 157]. Good [13, 157]. Good-bye [13]. Gotta [28]. Grandeur [67]. Grass [137]. Grill [213]. Guitar [166].

Halo [71]. Hand

[166, 64, 40, 209, 2, 185, 116, 80, 21, 97]. Hatter [127]. Hawaii [40]. Heade [7]. Heaven [71]. Heinrich [33]. Hell [25]. Her [211]. Hide [193]. High [32, 137, 71]. Highlights [16, 34, 51, 58, 91, 108, 125, 144, 160, 177, 196, 202, 110, 180, 94, 54, 128, 162, 18, 204, 199, 36, 61, 146]. Hispanic [146]. History [70, 12]. Home [46]. Horning [103]. House [101]. Houses [93].

ia [208]. Illusion [61]. I’m [193].

Imagemaker [184]. Images [199]. Imaging [97]. Imagining [153]. Imperial [199]. Impossible [164]. Increase [1]. India [199]. Indomitable [97]. Inscribing [214]. INSPIRATION [1]. Interest [115, 149, 183, 131, 63, 96, 213, 163, 20, 79, 165, 4, 38]. Introducing [4]. Iranian [24]. Island [13]. Isn’t [22]. Itself [168]. ItSteadyUpsy [18]. ItSteadyUpsy-daisy [18].

James [176]. Jeon [102]. Jeon-Siméon [102]. Jeu [8]. Jim [165]. Jitterbug [149, 163]. Job [164]. John [57]. Johnson [7]. Judge [173]. Jungle [165]. Just [121]. Justice [29].

Kills [151]. Kiwi [43]. Kudzu [169].

Label [106, 112]. Last

[106, 206, 131, 88, 158, 193, 48, 112, 13, 174, 71, 30, 21, 123, 141]. Leaf [133]. Leggers [100]. Legs [111]. Letters [17, 35, 52, 59, 92, 109, 126, 145, 161, 178, 197, 203, 192]. Libba [166]. Life [122]. Light [210, 24]. Lightning [210]. Like [136]. Liveliest [216]. Lofty [179]. Log [208]. Log-o-phil-ia [208]. Long [139]. Long-rifle [139]. Looking [122]. Love [90, 104, 169].

Machine [97]. Maclean [57]. Mad [127, 50]. Made [50]. Magma [23]. Making [5]. Malaria [151]. Mall [3, 55, 69, 78, 98, 133, 150, 184, 211]. Man [3, 192, 37]. Marlotte [152]. Martin [7]. Mason [87]. Mastery [102]. Me [47]. Mean [154]. Medical [97]. Mediterranean [15]. Men [50, 181]. Messages [75]. Michael [143]. Middleweight [66]. Mind [33]. Miniature [2]. Mom [36]. Monster [117]. Monumental [148]. Mountain [57]. Mountains [198]. Move [23]. Movement [204]. Moving [6, 78]. MRI [97]. Museum [216]. Mustard [104]. Myles [47]. Mystery [2].

N [57]. Nectar [72]. Nectar-feeding [72]. News [77, 95]. Newton [207]. Night [43]. Nightjars [121]. No [115, 106, 112, 82]. Noble [45]. Notable [186, 200]. Note [114, 73, 1, 164, 216, 148, 37, 76, 182, 130, 19, 113]. Notes [77, 132, 95, 39, 74, 167, 22]. Nouveau [170].

o [208]. Object [166, 64, 40, 209, 2, 185, 116, 80, 21, 97]. Off [97]. One [121, 64, 151]. Orange [117]. Order [113]. Orient [153]. Original [7]. Our [116]. Oz [105].

Page [106, 206, 88, 158, 193, 48, 112, 13, 174, 71, 30, 123, 141]. Pain [198]. Pants [48]. Paper [76]. Passionate [182]. Paume [8]. Pavord [50]. People [93, 211, 215]. Permafrost [22]. Petal [60]. Petrel [138]. Phage [167]. Phenomena [77, 132, 95, 39, 74, 167, 22]. phil [208]. Pi [82]. Pianos [53]. Picture [191]. Piece [82]. Pinhole [86]. Pixels [76]. Plains [80]. Play [42]. Poetic [119]. Point [86]. Points [115, 149, 183, 131, 63, 96, 213, 163, 20, 79, 165, 4, 38]. Portraits [211, 80]. Power [60]. Praise [53]. Presidential [185]. Presidents [187]. Princip [134]. Public [206]. Pursuit [29]. Pygmalion [188].

Quiet [102].

R [163]. R&B [149]. Race [209]. Rare [190]. Rarest [190]. Raven [33]. Raymond [97]. Reader [30]. Reading [130, 75]. Ready [9]. Real [143]. Reality [61]. Ream [140]. Recalls [110]. Recollecting [187]. Redefining [26]. Reindeer [215]. Remedy [209]. Remember [28]. Renaissance [139, 3]. Requiem [66]. Required [106, 112]. Rescue [129]. Responsibility [148]. Restoration [129]. Return [115, 167]. Review [176, 143]. Reviews [57, 90, 15, 33, 50, 210]. Revolution [97]. Revolutions [210]. Rhode [13]. Rhythm [204]. rifle [139]. Ringtails [136]. Rises [65]. Rivals [90]. Road [20]. Roadside [183]. Robot [135]. Robots [26]. Rock [101]. Rome [67]. Ruled [173]. Run [5, 169].

Santos [146]. Schoolhouse [131]. Sculptor [119]. Scythian [44]. Searching [134]. Second [41]. Seconds [151]. Secretary [114, 73, 1, 164, 216, 148, 37, 76, 182, 130, 19, 113]. Set [97]. Shod [218]. Shop [106, 112]. Shopping [48]. Short [48]. Signs [141]. Siméon [102]. Singing [94]. Sister [210]. Six [111]. Skies [114]. Sky [147]. Sloth [10]. Smithsonian [110, 180, 94, 54, 128, 162, 18, 204, 199, 36, 61, 146, 186, 200, 19]. Snow [218]. Something [135]. Soul [195, 146]. Sources [14, 31, 49, 56, 89, 107, 124, 142, 159, 175, 194, 201]. South [57]. Space [37]. Spanish [119]. Specialist [69]. Splendors [27]. Squaring [82]. Stellar [184]. Still [116]. Stink [71]. Storied [150]. Stories [174]. Stormy [138]. Story [57]. Stripe [36]. Stuff [6]. Supermodels [111].

SweetSo [154]. Swoopy [101].

Take [10]. Tale [39]. Talking [98]. Tea [38]. Tells [174]. Terrain [99]. Them [42]. There [116]. These [158]. Things [75]. Tickling [147]. Tigers [10]. Timeless [83]. Times [120, 141]. Tiny [154]. Too [158]. Topkapi [27]. Touch [99]. Trail [171]. Tribute [179]. True [57]. Try [121]. Tulip [50]. Turning [133]. Twigman [205]. Twit [88]. Two [81, 38].

Undertaking [45]. Unique [46].

Vanishing [40]. Vice [207]. View [86]. Vinnie [140]. Vision [119]. Vostok [122].

Walks [168]. Wallis [143]. Wanted [181, 63]. Wants [123]. War [110, 62, 217]. Warning [208]. Was [25, 67, 81, 116]. Waterlogging [156]. Wayne [90]. Weird [28]. Welcome [165]. West [143]. Wheels [81]. Who [42, 88, 173]. Whole [113]. Wild [143, 21]. Will [30]. Wind [41]. Without [97]. Witness [12]. Word [214]. World [5, 70]. Worlds [113]. Worm [132]. Wrecking [212]. Wright [15]. Wrong [191].

Yalta [12]. Yuppie [88].

References

Small:2000:NSIa

[1] Secretary Lawrence M. Small. A note from the Secretary: Increase, diffusion and … INSPIRATION. Smithsonian, 30(10):12–??, January 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jan00/small_
jan00.html.

Geracimos:2000:OHM

[2] Anne Geracimos. The object at hand: a mystery in miniature. Smithsonian, 30(10):20–??, January 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jan00/object_
jan00.html.

Kernan:2000:AMBa

[3] Michael Kernan. Around the Mall & beyond: a renaissance man. Smithsonian, 30(10):22–??, January 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jan00/mall_
jan00.html.

Webster:2000:IPI

[4] Donovan Webster. Introducing: Points of interest. Smithsonian, 30(10):26–??, January 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jan00/interest_
jan00.html.

Page:2000:MCR

[5] Jake Page. Making the chips that run the world. Smithsonian, 30 (10):36–??, January 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jan00/intel.

html. Photographs by Kay Chernush.

Clarke:2000:MBS

[6] Wendy Mitman Clarke. Moving big stuff. Smithsonian, 30(10):48– ??, January 2000. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues00/jan00/moving.html.

Schiff:2000:MJH

[7] Bennett Schiff. Martin Johnson Heade: An American original. Smithsonian, 30(10):62–??, January 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jan00/heade.
html.

Chelminski:2000:JPA

[8] Rudolph Chelminski. Jeu de paume, anyone? Smithsonian, 30(10):72– ??, January 2000. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues00/jan00/tennis.html. Pho-

tographs by Michael Freeman.

Hadingham:2000:RLA

[9] Evan Hadingham. Ready … aim … fire! Smithsonian, 30(10):78– ??, January 2000. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues00/jan00/trebuchets.html. Pho-

tographs by Patrick Ward.

Conover:2000:SBT

[10] Adele Conover. Sloth bears: They eat ants, but take on tigers. Smithsonian, 30(10):88–??, January 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/

                                        smithsonian/issues00/jan00/sloth.
                                        html.

Minnick:2000:BG

[11] Mimi Minnick. Breck girls. Smithsonian, 30(10):96–??, January 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/jan00/breck.
                                        html.

Wernick:2000:YWH

[12] Robert Wernick.  Yalta:  Wit-

ness to history. Smithsonian, 30 (10):100–??, January 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/jan00/yalta.
                                        html.

Jackson:2000:LPG

[13] Donald Dale Jackson. The last page: Good-bye, Rhode Island. Smithsonian, 30(10):120–??, January 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/jan00/last_
                                        jan00.html.

Anonymous:2000:ASa

[14] Anonymous. Additional sources. Smithsonian, 30(10):??, January 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/jan00/additional_
                                        jan00.html.

Anonymous:2000:BRMa

[15] Anonymous. Book reviews: A Mediterranean Feast, by Clifford A. Wright.

Smithsonian, 30(10):??, January 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jan00/bookreview_
jan00.html.

Anonymous:2000:HCa

[16] Anonymous. Highlights calendar. Smithsonian, 30(10):??, January 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jan00/highcal_
jan00.html.

Anonymous:2000:LEa

[17] Anonymous. Letters to the Editor. Smithsonian, 30(10):??, January 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jan00/letters_
jan00.html.

Anonymous:2000:SHE

[18] Anonymous. Smithsonian highlights: Easy does it… steady… upsy-daisy. Smithsonian, 30(10):??, January 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jan00/highopener_
jan00.html.

Small:2000:NSS

[19] Secretary Lawrence M. Small. A note from the Secretary: a Smithsonian for this century. Smithsonian, 30(11):14–??, February 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/small_
feb00.html.

Watkins:2000:PIE

[20] T. H. Watkins. Points of interest: The end of the road. Smithsonian, 30(11):22–??, February 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/interest_
feb00.html.

Shell:2000:OHL

[21] Hanna Rose Shell. The object at hand: Last of the wild buffalo. Smithsonian, 30(11):26–??, February 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/object_
feb00.html.

Wuethrich:2000:PCN

[22] Bernice Wuethrich. Phenomena, comment & notes: When permafrost isn’t. Smithsonian, 30 (11):32–??, February 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/phenom_
feb00.html.

Wheeler:2000:WMM

[23] Mark Wheeler. When magma’s on the move. Smithsonian, 30 (11):40–??, February 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/magma.

html. Photographs by Eric Sander.

Moffett:2000:CLI

[24] Mark W. Moffett. Casting light on Iranian deserts. Smithsonian, 30(11):50–??, February 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/iran.

html. Photographs by Mark W. Moffett.

Bond:2000:FHW

[25] Constance Bond. A fury from Hell — or was he? Smithsonian, 30 (11):62–??, February 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/blackbeard.
html.

Trachtman:2000:RR

[26] Paul Trachtman. Redefining robots. Smithsonian, 30(11):96–??, February 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/feb00/

robots.html. Photographs by Peter Menzel.

Meisler:2000:ST

[27] Stanley Meisler. Splendors of Topkapi. Smithsonian, 30(11):114–??, February 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/feb00/
topkapi.html.

Watson:2000:YGR

[28] Bruce Watson. “you gotta remember, eels are weird.”. Smithsonian, 30(11):124–??, February 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/eels.
html.

Wolkomir:2000:DDP

[29] Richard Wolkomir and Joyce Wolkomir. Day by day, in pursuit of justice. Smithsonian, 30(11):134–??, February 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/justice.

html. Illustrations by Carolyne D. Landon.

Rodgers:2000:LPB

[30] Mary Augusta Rodgers. The last page: For a bedtime reader, not any book will do. Smithsonian, 30 (11):156–??, February 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/last_
feb00.html.

Anonymous:2000:ASb

[31] Anonymous. Additional sources. Smithsonian, 30(11):??, February 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/additional_
feb00.html.

Anonymous:2000:BH

[32] Anonymous. Beasts on high. Smithsonian, 30(11):??, February 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/gargoyles.

html. Photographs by Robert Llewellyn.

Anonymous:2000:BRMb

[33] Anonymous. Book reviews: Mind of the Raven, by Bernd Heinrich. Smithsonian, 30(11):??, February 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/bookreview_
feb00.html.

Anonymous:2000:HCb

[34] Anonymous. Highlights calendar. Smithsonian, 30(11):??, February 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/highcal_
feb00.html.

Anonymous:2000:LEb

[35] Anonymous. Letters to the Editor. Smithsonian, 30(11):??, February 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/letters_
feb00.html.

Anonymous:2000:SHM

[36] Anonymous. Smithsonian highlights: Mom earns another stripe. Smithsonian, 30(11):??, February 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/feb00/highopener_
feb00.html.

Small:2000:NSN

[37] Secretary Lawrence M. Small. A note from the Secretary: The new man at Air and Space. Smithsonian, 30(12):14–??, March 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/small_
mar00.html.

Webster:2000:PIT

[38] Donovan Webster. Points of interest: Two for tea. Smithsonian, 30(12):24–??, March 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/interest_
mar00.html.

Park:2000:PCN

[39] Edwards Park. Phenomena, comment & notes: a tale of fatal feuds and futile forensics. Smithsonian, 30(12):32–??, March 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/phenom_
mar00.html.

Conover:2000:OHH

[40] Adele Conover. The object at hand: Hawaii’s vanishing birds. Smithsonian, 30(12):38–??, March 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/object_
mar00.html.

Chiles:2000:SW

[41] James Chiles. A second wind. Smithsonian, 30(12):50–??, March 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/wind.

html. Photographs by Layne Kennedy.

Chelminski:2000:AWP

[42] Rudolph Chelminski. And the artists who play them. Smithsonian, 30(12):64–??, March 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/pianists.

html. Illustrations by William Bramhall.

Grzelewski:2000:NBK

[43] Derek Grzelewski. Night belongs to the kiwi. Smithsonian, 30(12):76–??, March 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/mar00/
kiwi.html.

Stewart:2000:SG

[44] Doug Stewart. Scythian gold. Smithsonian, 30(12):88–??, March 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/gold.
html.

Taylor:2000:NAU

[45] David A. Taylor. A noble and absurd undertaking. Smithsonian, 30(12):100–??, March 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/wpa.
html.

Watson:2000:UHC

[46] Bruce Watson. A unique home for cutting-edge art. Smithsonian, 30(12):114–??, March 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/moca.

html. Photographs by Frederick C. Charles.

Standish:2000:MM

[47] David Standish. Myles and me. Smithsonian, 30(12):126–??, March 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/standish.

html. Illustrations by Rob Day.

Gup:2000:LPH

[48] Ted Gup. The last page: He’s shopping for pants but coming up short. Smithsonian, 30(12):148–??, March 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/last_
mar00.html.

Anonymous:2000:ASc

[49] Anonymous. Additional sources. Smithsonian, 30(12):??, March 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/additional_
mar00.html.

Anonymous:2000:BRT

[50] Anonymous. Book reviews: The Tulip: The Flower that Has Made Men Mad, by Anna Pavord. Smithsonian, 30(12):??, March 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/bookreview_
mar00.html.

Anonymous:2000:HCc

[51] Anonymous. Highlights calendar. Smithsonian, 30(12):??, March 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/highcal_
mar00.html.

Anonymous:2000:LEc

[52] Anonymous. Letters to the Editor. Smithsonian, 30(12):??, March 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/letters_
mar00.html.

Anonymous:2000:PPa

[53] Anonymous. In praise of pianos. . .. Smithsonian, 30(12):??, March 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/pianos.

html. Photographs by Robert Lautman.

Anonymous:2000:SHDa

[54] Anonymous. Smithsonian highlights: Driven to explore America. Smithsonian, 30(12):??, March 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/highopener_
mar00.html.

Kernan:2000:AMBb

[55] Michael Kernan. Around the Mall & beyond: Bonding through books. Smithsonian, 30(12):??, March 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/mar00/mall_
mar00.html.

Anonymous:2000:ASd

[56] Anonymous. Additional sources. Smithsonian, 31(1):??, April 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/additional_
apr00.html.

Anonymous:2000:BRFa

[57] Anonymous. Book reviews: Fire on the Mountain: The True Story of the South Canyon Fire, by John N. Maclean. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/bookreview_
apr00.html.

Anonymous:2000:HCd

[58] Anonymous. Highlights calendar. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/highcal_
apr00.html.

Anonymous:2000:LEd

[59] Anonymous. Letters to the Editor. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/letters_
apr00.html.

Anonymous:2000:PPb

[60] Anonymous. Petal power. Smithsonian, 31(1):??, April 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/flowers.

html. Photographs by Amy Lamb.

Anonymous:2000:SHR

[61] Anonymous. Smithsonian highlights: The reality of illusion. Smithsonian, 31(1):??, April 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/highopener_
apr00.html.

Chettle:2000:WWC

[62] John H. Chettle. When war called, Davis answered. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/apr00/
davis.html.

Conniff:2000:PIW

[63] Richard Conniff. Points of interest: Wanted, dead or alive. Smithsonian, 31(1):??, April 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/interest_
apr00.html.

Conover:2000:OHB

[64] Adele Conover. The object at hand: The biggest one that didn’t get away. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/object_
apr00.html.

Covington:2000:BRA

[65] Richard Covington. Beirut rises from the ashes. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/apr00/

beirut.html. Photographs by Robert Wallis.

Doherty:2000:RM

[66] Jim Doherty. Requiem for a middleweight. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/apr00/
boxing.html.

Dudar:2000:GWR

[67] Helen Dudar. The grandeur that was rome. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/rome.
html.

Gup:2000:ED

[68] Ted Gup. Empire of the dead. Smithsonian, 31(1):??, April 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/catacombs.
html.

Kernan:2000:AMBc

[69] Michael Kernan. Around the Mall & beyond: Bone specialist on call. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/mall_
apr00.html.

Kurlansky:2000:BHW

[70] Mark Kurlansky. The Basque history of the World. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Lemen:2000:LPE

[71] June Lemen. The last page: Earning a halo can stink to high heaven. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/last_
apr00.html.

Schwartz:2000:BBE

[72] David M. Schwartz. Birds, bees and even nectar-feeding bats do it. Smithsonian, 31(1):??, April 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/

smithsonian/issues00/apr00/pollen.
html.

Small:2000:NSG

[73] Secretary Lawrence M. Small. A note from the Secretary: The gem of the collection. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/apr00/
small_apr00.html.

Trefil:2000:PCN

[74] James Trefil. Phenomena, comment & notes: Evidence for a flood. Smithsonian, 31(1):??, April 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/apr00/phenom_
apr00.html.

Wolkomir:2000:RME

[75] Richard Wolkomir and Joyce Wolkomir. Reading the messages in everyday things. Smithsonian, 31(1):??, April 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/apr00/
stilgoe.html.

Small:2000:NSPa

[76] Secretary Lawrence M. Small. A note from the Secretary: From paper to pixels. Smithsonian, 31(2):18–??, May 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/may00/
small_may00.html.

Anonymous:2000:PCN

[77] Anonymous. Phenomena, comment & notes: a celestial news bureau. Smithsonian, 31(2):28–??, May 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:2000:AMBd

[78] Michael Kernan. Around the Mall & beyond: Art’s moving experience. Smithsonian, 31(2):30–??, May 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/mall_
may00.html.

Watson:2000:PIB

[79] Bruce Watson. Points of interest: Bound for glory. Smithsonian, 31(2):36–??, May 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/interest_
may00.html.

Park:2000:OHP

[80] Edwards Park. The object at hand: Portraits on the plains. Smithsonian, 31(2):42–??, May 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/object_
may00.html.

Parfit:2000:CWT

[81] Michael Parfit. “the corn was two feet below the wheels”. Smithsonian, 31(2):54–??, May 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/cropdusting.

html. Photographs by Chad Slattery.

Watson:2000:SCN

[82] Bruce Watson. Squaring the circle is no piece of pi. Smithsonian, 31(2):70–??, May 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/pi.html.

Illustrations by Jeff Seaver.

Bolz:2000:TAC

[83] Diane M. Bolz. The timeless art of crafts. Smithsonian, 31(2):86–??, May 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/may00/

crafts.html. Photographs by Bill Ballenberg.

Grzelewski:2000:BAF

[84] Derek Grzelewski. Batty about flying foxes. Smithsonian, 31(2):100–??, May 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/may00/

bats.html. Photographs by Theo Allofs.

Page:2000:BB

[85] Jake Page. Beauty and the beasts. Smithsonian, 31(2):110–??, May 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/makeup.

html. Photographs by Eric Sander.

Stewart:2000:PPV

[86] Doug Stewart. The pinhole point of view. Smithsonian, 31(2):124–??, May 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/may00/
pinhole.html.

Schwartz:2000:GMF

[87] Stephan A. Schwartz. George Mason: Forgotten founder. Smithsonian, 31(2):142–??, May 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/mason.
html.

Conniff:2000:LPW

[88] Richard Conniff. The last page: Who’s a yuppie twit? Smithsonian, 31(2):160–??, May 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/last_
may00.html.

Anonymous:2000:ASe

[89] Anonymous. Additional sources. Smithsonian, 31(2):??, May 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/additional_
may00.html.

Anonymous:2000:BRL

[90] Anonymous. Book reviews: For the Love of It: Amateuring and Its Rivals, by Wayne Booth. Smithsonian, 31(2):??, May 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/bookreview_
may00.html.

Anonymous:2000:HCe

[91] Anonymous. Highlights calendar. Smithsonian, 31(2):??, May 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/highcal_
may00.html.

Anonymous:2000:LEe

[92] Anonymous. Letters to the Editor. Smithsonian, 31(2):??, May 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/letters_
may00.html.

Anonymous:2000:PGH

[93] Anonymous. People in glass houses. . .. Smithsonian, 31(2):??, May 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/glass.

html. Photographs by Michael Freeman.

Anonymous:2000:SHC

[94] Anonymous. Smithsonian highlights: “I commenced to singing”. Smithsonian, 31(2):??, May 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/highopener_
may00.html.

Jablow:2000:PCN

[95] Valerie Jablow. Phenomena, comment & notes: a celestial news bureau. Smithsonian, 31(2):??, May 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/may00/phenom_
may00.html.

Morse:2000:PIC

[96] Minna Morse. Points of interest: Camping in concert. Smithsonian, 31(3):16–??, June 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jun00/interest_
jun00.html.

Wakefield:2000:OHI

[97] Julie Wakefield. The object at hand: The “indomitable” MRI: Raymond Damadian’s medical imaging machine set off a revolution — but not without controversy. Smithsonian, 31(3):22–??, June 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jun00/object_
jun00.html.

Kernan:2000:AMBe

[98] Michael Kernan. Around the Mall & beyond: The talking drums. Smithsonian, 31(3):27–??, June 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jun00/mall_
jun00.html.

Wolkomir:2000:CTT

[99] Richard Wolkomir. Charting the terrain of touch. Smithsonian, 31(3): 38–??, June 2000. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

                                        smithsonianmag.si.edu/smithsonian/
                                        issues00/jun00/touch.html. Pho-

tographs by Richard Schultz.

Hornik:2000:AFB

[100] Susan Hornik. Afloat with fly boats and leggers. Smithsonian, 31(3):50–??, June 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues00/jun00/

canals.html. Photographs by Michael Freeman.

Covington:2000:SFF

[101] Richard Covington. A swoopy, funky fun house of rock. Smithsonian, 31(3):68–??, June 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/jun00/music.

html. Photographs by Steve Gottlieb.

Tuchman:2000:QMJ

[102] Phyllis Tuchman. The quiet mastery of Jeon-Siméon Chardin. Smithsonian, 31(3):80–??, June 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/jun00/chardin.
                                        html.

Kemper:2000:H

[103] Steve Kemper. Horning in? Smithsonian, 31(3):90–??, June 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jun00/sheep.
html.

Harriss:2000:LM

[104] Joseph A. Harriss. For the love of mustard. Smithsonian, 31(3):102–??, June 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/jun00/

mustard.html. Photographs by Brooks Walker.

Watson:2000:AAO

[105] Bruce Watson. The amazing author of Oz. Smithsonian, 31(3):112–??, June 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/jun00/
oz.html.

Anonymous:2000:LPS

[106] Anonymous. The last page: Shop for this label: “no assembly required.”. Smithsonian, 31(3):124–??, June 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:2000:ASf

[107] Anonymous. Additional sources. Smithsonian, 31(3):??, June 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jun00/additional_
jun00.html.

Anonymous:2000:HCf

[108] Anonymous. Highlights calendar.

Smithsonian, 31(3):??, June 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jun00/highcal_
jun00.html.

Anonymous:2000:LEf

[109] Anonymous. Letters to the Editor. Smithsonian, 31(3):??, June 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jun00/letters_
jun00.html.

Anonymous:2000:SHA

[110] Anonymous. Smithsonian highlights: Art recalls a forgotten war. Smithsonian, 31(3):??, June 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jun00/highopener_
jun00.html.

Anonymous:2000:SSL

[111] Anonymous. Supermodels with six legs. Smithsonian, 31(3):??, June 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jun00/bugs.

html. Photographs by Gilles Mermet.

Heitman:2000:LPS

[112] Danny Heitman. The last page: Shop for this label: “no assembly required”. Smithsonian, 31(3):??, June 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jun00/last_
jun00.html.

Small:2000:NSW

[113] Secretary Lawrence M. Small. A note from the Secretary: Whole worlds to order. Smithsonian, 31(3):??, June 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jun00/small_
jun00.html.

Small:2000:NSE

[114] Secretary Lawrence M. Small. A note from the Secretary: Evidence from the skies. Smithsonian, 31(4):12–??, July 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/jul00/
small_jul00.html.

Allison:2000:PIN

[115] Sue Allison. Points of interest: No return address. Smithsonian, 31(4): 18–??, July 2000. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues00/jul00/interest_jul00.html.

Park:2000:OHO

[116] Edwards Park. The object at hand:

Our flag was still there. Smithsonian, 31(4):22–??, July 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jul00/object_
jul00.html.

Kiester:2000:BOM

[117] Edwin Kiester, Jr. Battling the orange monster. Smithsonian, 31(4):32–??, July 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/jul00/

fire.html. Photographs by Ed Kashi.

Moore:2000:CAF

[118] Lucinda Moore. Capturing America’s fight for freedom. Smithsonian, 31(4):44–??, July 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jul00/patriot.
html.

Meisler:2000:PVS

[119] Stanley Meisler. The poetic vision of Spanish sculptor Eduardo Chillida. Smithsonian, 31(4):56–??, July 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jul00/chillida.
html.

Shepherd:2000:FTT

[120] Virginia Shepherd. Down on this farm the times they are A-changin’. Smithsonian, 31(4):64–??, July 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jul00/farm.

html. Photographs by Lynda Richardson.

Burt:2000:NEJ

[121] William Burt. Nightjars are everywhere, but just try finding one. Smithsonian, 31(4):74–??, July 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jul00/nightjars.
html.

Stone:2000:VLL

[122] Richard Stone. Vostok: Looking for life beneath an Antarctic glacier. Smithsonian, 31(4):92–??, July 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jul00/vostok.
html.

Watson:2000:LPE

[123] Bruce Watson. The last page: Every dog wants to have its day in court. Smithsonian, 31(4):104–??, July 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jul00/last_
jul00.html.

Anonymous:2000:ASg

[124] Anonymous. Additional sources. Smithsonian, 31(4):??, July 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jul00/additional_
jul00.html.

Anonymous:2000:HCg

[125] Anonymous. Highlights calendar.

Smithsonian, 31(4):??, July 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/jul00/highcal_
jul00.html.

Anonymous:2000:LEg

[126] Anonymous. Letters to the Editor. Smithsonian, 31(4):??, July 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/jul00/letters_
                                        jul00.html.

Anonymous:2000:MH

[127] Anonymous. The Mad Hatter. Smithsonian, 31(4):??, July 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/jul00/hats.
                                        html.

Anonymous:2000:SHDb

[128] Anonymous. Smithsonian highlights: It’s a drum. Smithsonian, 31(4):??, July 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/jul00/highopener_
                                        jul00.html.

Holmes:2000:RR

[129] Marian Smith Holmes. Rescue and restoration. Smithsonian, 31(4):??, July 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues00/jul00/
                                        hbcus.html.

Small:2000:NSR

[130] Secretary Lawrence M. Small. A note from the Secretary: Reading the bones. Smithsonian, 31(5):16–??, August 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/aug00/small_
aug00.html.

Chelminski:2000:PIL

[131] Rudolph Chelminski. Points of interest: The last schoolhouse. Smithsonian, 31(5):22–??, August 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/aug00/interest_
aug00.html.

Conover:2000:PCN

[132] Adele Conover. Phenomena, comment & notes: Foreign worm alert. Smithsonian, 31(5):29–??, August 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/aug00/phenom_
aug00.html.

Kernan:2000:AMBf

[133] Michael Kernan. Around the Mall & beyond: Turning a new leaf. Smithsonian, 31(5):32–??, August 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/aug00/mall_
aug00.html.

DeVoss:2000:SGP

[134] David DeVoss. Searching for Gavrilo Princip. Smithsonian, 31(5):42– ??, August 2000. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues00/aug00/princip.html. Pho-

tographs by Manfred Horvath.

Whynott:2000:SFA

[135] Douglas Whynott. Something’s fishy about this robot. Smithsonian, 31(5):54–??, August 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/aug00/robofish.

html. Photographs by Peter Menzel.

Gilbert:2000:RLA

[136] Bil Gilbert. Ringtails like to be appreciated. Smithsonian, 31(5):64–??, August 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues00/aug00/
                                        ringtails.html.

Kemper:2000:HG

[137] Steve Kemper. High on grass. Smithsonian, 31(5):72–??, August 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/aug00/park.

html. Photographs by Theo Westenberger.

Ferris:2000:SPA

[138] Scott R. Ferris. “the stormy petrel of American art.”. Smithsonian, 31(5):78–??, August 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/aug00/kent.
                                        html.

Conaway:2000:RLR

[139] James Conaway. Renaissance of the long-rifle. Smithsonian, 31(5):

96–??, August 2000. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues00/aug00/longrifle.html. Pho-

tographs by Walter P. Calahan.

Jacob:2000:VR

[140] Kathryn Allamong Jacob. Vinnie Ream.

Smithsonian, 31(5):104–??, August 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/aug00/
ream.html.

White:2000:LPS

[141] Robb White. The last page: Signs of the times. Smithsonian, 31 (5):120–??, August 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/aug00/last_
aug00.html.

Anonymous:2000:ASh

[142] Anonymous. Additional sources. Smithsonian, 31(5):??, August 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/aug00/additional_
aug00.html.

Anonymous:2000:BRR

[143] Anonymous. Book review: The Real Wild West, by Michael Wallis. Smithsonian, 31(5):??, August 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/aug00/bookreview_
aug00.html.

Anonymous:2000:HCh

[144] Anonymous. Highlights calendar. Smithsonian, 31(5):??, August 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/aug00/highcal_
                                        aug00.html.

Anonymous:2000:LEh

[145] Anonymous. Letters to the Editor. Smithsonian, 31(5):??, August 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/aug00/letters_
                                        aug00.html.

Anonymous:2000:SHS

[146] Anonymous. Smithsonian highlights: The soul of the Hispanic Santos. Smithsonian, 31(5):??, August 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/aug00/highopener_
                                        aug00.html.

Anonymous:2000:TS

[147] Anonymous. Tickling the sky. Smithsonian, 31(5):??, August 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/aug00/gazit.

html. Photographs by Eric Sander & Doron Gazit.

Small:2000:NSM

[148] Secretary Lawrence M. Small. A note from the Secretary: a monumental responsibility. Smithsonian, 31(6):12–??, September 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/sep00/small_
sep00.html.

Anonymous:2000:PIJ

[149] Anonymous. Points of interest: “the jitterbug met R&B”. Smithsonian, 31(6): 16–??, September 2000. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Kernan:2000:AMBg

[150] Michael Kernan. Around the Mall & beyond: a storied gallery. Smithsonian, 31(6):22–??, September 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/sep00/mall_
sep00.html.

Webster:2000:MKO

[151] Donovan Webster. Malaria kills one child every 30 seconds. Smithsonian, 31(6):32–??, September 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/sep00/malaria.

html. Photographs by Michael Freeman.

Chelminski:2000:ABM

[152] Rudolph Chelminski. An American in Bourron-Marlotte. Smithsonian, 31(6):46–??, September 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/sep00/bourron.

html. Illustrations by Robert Gantt

Steele.

Stewart:2000:IO

[153] Doug Stewart. Imagining the Orient. Smithsonian, 31(6):60–??, September 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues00/sep00/
                                        orient.html.

Conniff:2000:TSM

[154] Richard Conniff. So tiny, so sweet… so mean. Smithsonian, 31 (6):72–??, September 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/sep00/hummingbirds.
                                        html.

Millar:2000:DA

[155] Heather Millar. Dancing on air. Smithsonian, 31(6):86–??, September 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/sep00/bandaloop.

html. Photographs by Peter McBride.

Wakefield:2000:W

[156] Julie Wakefield. Waterlogging. Smithsonian, 31(6):94–??, September 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/sep00/waterlogging.

html. Photographs by Layne Kennedy.

Stiles:2000:GG

[157] T. J. Stiles. As good as gold? Smithsonian, 31(6):106–??, September 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/sep00/gold.
html.

Dumas:2000:LPF

[158] Gerald Dumas. The last page: a fellow can’t be too careful these days. Smithsonian, 31(6):120–??, September 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/sep00/last_
sep00.html.

Anonymous:2000:ASi

[159] Anonymous. Additional sources. Smithsonian, 31(6):??, September 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/sep00/additional_
sep00.html.

Anonymous:2000:HCi

[160] Anonymous. Highlights calendar.

Smithsonian, 31(6):??, September 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/sep00/highcal_
sep00.html.

Anonymous:2000:LEi

[161] Anonymous. Letters to the Editor.

Smithsonian, 31(6):??, September 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/sep00/letters_
sep00.html.

Anonymous:2000:SHDc

[162] Anonymous. Smithsonian highlights: Diamond days. Smithsonian, 31(6):??, September 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/sep00/highopener_
                                        sep00.html.

Nickens:2000:PIJ

[163] T. Edward Nickens. Points of interest: “the jitterbug met R”. Smithsonian, 31(6):??, September 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/sep00/interest_
                                        sep00.html.

Small:2000:NSIb

[164] Secretary Lawrence M. Small. A note from the Secretary: The impossible job. Smithsonian, 31(7):16–??, October 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues00/oct00/
                                        small_oct00.html.

Watson:2000:PIW

[165] Bruce Watson. Points of interest: Welcome to Jungle Jim’s. Smithsonian, 31(7):26–??, October 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/oct00/interest_
                                        oct00.html.

Carley:2000:OHL

[166] Marika Carley. The object at hand: Libba Cotten’s guitar. Smithsonian, 31(7):32–??, October 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/oct00/object_
oct00.html.

Wakefield:2000:PCN

[167] Julie Wakefield. Phenomena, comment & notes: The return of the phage. Smithsonian, 31(7):43–??, October 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/oct00/
phenom_oct00.html.

Bergman:2000:CWI

[168] Charles Bergman. The cat that walks by itself. Smithsonian, 31 (7):54–??, October 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/oct00/jaguar.
html.

Stewart:2000:KLI

[169] Doug Stewart.  Kudzu:  Love

it — or run. Smithsonian, 31 (7):64–??, October 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/oct00/kudzu.

html. Photographs by Kathryn Kolb.

Meisler:2000:AN

[170] Stanley Meisler. Art nouveau. Smithsonian, 31(7):74–??, October 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/

                                        smithsonian/issues00/oct00/artnouveau.
                                        html.

McCaig:2000:BT

[171] Donald McCaig. The Bozeman Trail. Smithsonian, 31(7):88–??, October 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues00/oct00/

bozeman.html. Illustrations by Francis Livingston.

Wolkomir:2000:YWY

[172] Richard Wolkomir and Joyce Wolkomir. You are what you buy. Smithsonian, 31(7):102–??, October 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/oct00/ads.
                                        html.

Watson:2000:JWR

[173] Bruce Watson. The judge who ruled baseball. Smithsonian, 31 (7):120–??, October 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/oct00/landis.
                                        html.

Kaplan:2000:LPB

[174] Janice L. Kaplan. The last page: The book that tells stories about you. Smithsonian, 31(7):140–??, October 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/oct00/last_
                                        oct00.html.

Anonymous:2000:ASj

[175] Anonymous. Additional sources. Smithsonian, 31(7):??, October 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/oct00/additional_
oct00.html.

Anonymous:2000:BRFb

[176] Anonymous. Book review: Faster, by James Gleick. Smithsonian, 31 (7):??, October 2000. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues00/oct00/bookreview_oct00.
html.

Anonymous:2000:HCj

[177] Anonymous. Highlights calendar.

Smithsonian, 31(7):??, October 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/oct00/highcal_
oct00.html.

Anonymous:2000:LEj

[178] Anonymous. Letters to the Editor.

Smithsonian, 31(7):??, October 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/oct00/letters_
oct00.html.

Anonymous:2000:LTB

[179] Anonymous. A lofty tribute to barns.

Smithsonian, 31(7):??, October 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/oct00/barns.

html. Photographs by David Plowden.

Anonymous:2000:SHB

[180] Anonymous. Smithsonian highlights: Building a better dinosaur. Smithsonian, 31(7):??, October 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/oct00/highopener_
oct00.html.

Anonymous:2000:WBM

[181] Anonymous. Wanted: Big men. Smithsonian, 31(7):??, October 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/oct00/mufflermen.
html.

Small:2000:NSPb

[182] Secretary Lawrence M. Small. A note from the Secretary: a passionate collector. Smithsonian, 31 (8):18–??, November 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/small_
nov00.html.

Carter:2000:PIR

[183] W. Hodding Carter. Points of interest: Roadside attractions. Smithsonian, 31(8):28–??, November 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/interest_
nov00.html.

Kernan:2000:AMBh

[184] Michael Kernan. Around the Mall & beyond: a stellar imagemaker. Smithsonian, 31(8):34–??, November 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/mall_
nov00.html.

Jarlow:2000:OHP

[185] Valerie Jarlow. The object at hand:

Presidential designs. Smithsonian, 31(8):46–??, November 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/object_
nov00.html.

Anonymous:2000:SNB

[186] Anonymous. Smithsonian’s notable books for children, 2000. Smithsonian, 31(8):58–??, November 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

Anonymous:2000:RP

[187] Anonymous. Recollecting the presidents. Smithsonian, 31(8):74–??, November 2000. CODEN SM-SNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL http://www.

smithsonianmag.si.edu/smithsonian/
issues00/nov00/presidents.html.

Harriss:2000:PAG

[188] Joseph A. Harriss. The Pygmalion of the avant-garde. Smithsonian, 31(8):88–??, November 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/guillaume.
html.

Ola:2000:BB

[189] Per Ola and Emily D’Aulaire. Baking up a business. Smithsonian, 31 (8):108–??, November 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/flour.
html.

Weidensaul:2000:RR

[190] Scott Weidensaul. The rarest of the rare. Smithsonian, 31(8):118–??, November 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/nov00/
conservation.html. Photographs by

Cameron Davidson.

Watson:2000:WWP

[191] Bruce Watson. What’s wrong with this picture? Smithsonian, 31 (8):130–??, November 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/pageant.

html. Photographs by Seth Joel.

Newman:2000:ACM

[192] Lisa Newman. Andrew Carroll: Man of letters. Smithsonian, 31 (8):152–??, November 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/carroll.

html. Photographs by Dennis Drenner.

Dumas:2000:LPH

[193] Gerald Dumas. The last page: Hide the catalogs — I’m out of control. Smithsonian, 31(8):172–??, November 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/last_
nov00.html.

Anonymous:2000:ASk

[194] Anonymous. Additional sources. Smithsonian, 31(8):??, November 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/additional_
nov00.html.

Anonymous:2000:FBS

[195] Anonymous. Farewell, bright soul.

Smithsonian, 31(8):??, November 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/gravestones.

html. Photographs by Gregory Thorp.

Anonymous:2000:HCk

[196] Anonymous. Highlights calendar.

Smithsonian, 31(8):??, November 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/highcal_
nov00.html.

Anonymous:2000:LEk

[197] Anonymous. Letters to the Editor.

Smithsonian, 31(8):??, November 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/letters_
nov00.html.

Anonymous:2000:MP

[198] Anonymous. Mountains of pain. Smithsonian, 31(8):??, November 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/bolivia.

html. Photographs by Jean-Claude Wicky.

Anonymous:2000:SHI

[199] Anonymous. Smithsonian highlights: Images from Imperial India. Smithsonian, 31(8):??, November 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/highopener_
nov00.html.

Burke:2000:SNB

[200] Kathleen Burke. Smithsonian’s Notable Books for Children 2000. Smithsonian, 31(8):??, November 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/nov00/bookreview_
nov00.html.

Anonymous:2000:ASl

[201] Anonymous. Additional sources. Smithsonian, 31(9):??, December 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/dec00/additional_
dec00.html.

Anonymous:2000:HCl

[202] Anonymous. Highlights calendar.

Smithsonian, 31(9):??, December 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/dec00/highcal_
dec00.html.

Anonymous:2000:LEl

[203] Anonymous. Letters to the Editor.

Smithsonian, 31(9):??, December 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/dec00/letters_
dec00.html.

Anonymous:2000:SHG

[204] Anonymous. Smithsonian highlights: a gift of rhythm and movement. Smithsonian, 31(9):??, December 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/dec00/highopener_
dec00.html.

Anonymous:2000:TC

[205] Anonymous. The twigman cometh.

Smithsonian, 31(9):??, December 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/dec00/dougherty.
html.

Black:2000:LPP

[206] Kent Black. The last page: The public bath. Smithsonian, 31(9):??, December 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/dec00/
last_dec00.html.

Carrell:2000:NV

[207] Jennifer Lee Carrell. Newton’s vice. Smithsonian, 31(9):??, December 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/dec00/newton.
html.

Chelminski:2000:WLP

[208] Rudolph Chelminski. Warning: Logo-phil-ia is addictive. Smithsonian, 31(9):??, December 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/dec00/wordaday.

html. Photographs by Mark Tuschman.

Crowley:2000:OHR

[209] Carolyn H. Crowley. The object at hand: Race for a remedy. Smithsonian, 31(9):??, December 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/dec00/object_
dec00.html.

Dunn:2000:BRS

[210] Susan Dunn. Book reviews: Sister revolutions: French lightning, American light. Smithsonian, 31(9):??, December 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

URL http://www.smithsonianmag.
si.edu/smithsonian/issues00/dec00/
bookreview_dec00.html.

Kernan:2000:AMBi

[211] Michael Kernan. Around the Mall & beyond: Portraits of her people. Smithsonian, 31(9):??, December 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/dec00/mall_
dec00.html.

Leibowitz:2000:UWB

[212] Ed Leibowitz. Out from under the wrecking ball. Smithsonian, 31(9):??, December 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/dec00/modcom.

html. Photographs by Marshall La-Plante.

Nickens:2000:PIC

[213] T. Edward Nickens. Points of interest: The Cypress Grill. Smithsonian, 31(9):??, December 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/dec00/interest_
dec00.html.

Ola:2000:IW

[214] Per Ola and Emily D’Aulaire. Inscribing the word. Smithsonian, 31(9):??, December 2000. CO-DEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

http://www.smithsonianmag.si.edu/
smithsonian/issues00/dec00/bible.

html. Photographs by Michael Freeman.

Ross:2000:PR

[215] John F. Ross. People of the reindeer. Smithsonian, 31(9):??, December 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/dec00/khanty.

html. Photographs by Scott Warren.

Small:2000:NSL

[216] Secretary Lawrence M. Small. A note from the Secretary: The liveliest museum. Smithsonian, 31(9):??, December 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic).

                                        URL http://www.smithsonianmag.
                                        si.edu/smithsonian/issues00/dec00/
                                        small_dec00.html.

Watson:2000:DW

[217] Bruce Watson. The dogs of war. Smithsonian, 31(9):??, December 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/dec00/wardogs.
                                        html.

Wolkomir:2000:SS

[218] Richard Wolkomir. Shod for snow. Smithsonian, 31(9):??, December 2000. CODEN SMSNA5. ISSN 0037-7333 (print), 1930-5508 (electronic). URL

                                        http://www.smithsonianmag.si.edu/
                                        smithsonian/issues00/dec00/snowshoes.

html. Photographs by David McLain.